is a graph of V versus I. V (V) 50 ºt x (cm) Ź 3
Part A R 4 of Select the correct graph of E, versus a. E.(V/m) 5000 x(cm) 5000 E (V/m) 5000 *(cm) 000
E. (V/m) S000 - *(cm) -S000 E. (V/m) 5000 -x(cm) 5000
Answer #1

20 2 3 (30) alcm). from ato to t= 2 ACCm) ↑ (2,50) slope of line = 50-(50) m = (1) sorlem equation of line -506-50) n= mate =E = - du y = - d (-50x1150) - 50 v/m = 5000 v/m. da therefore E=S-5000 ulm on 1 + 5000 v/m 02x42cm 2<x<3um 5000 + -5000